calculate the percent by mass of solute at 12.0 g of potassium carbonate in 100.0 g of water.


1
Expert's answer
2021-01-04T03:46:28-0500

We are to find the percentage weight of potassium carbonate (The solute) in water.

%w/w= weight of solute ÷weight of solvent x 100%

%w/w= 12g/100g x100%

%w/w = 12%(w/w)